# Beyond Reactivity: Measuring Proactive Problem Solving in LLM Agents

### v2

@@ Abstract
LLM-based agents are increasingly moving towards proactivity: rather than awaiting instruction, they exercise agency to anticipate user needs and solve them autonomously. However, evaluating proactivity is challenging; current benchmarks are constrained to localized context, limiting their ability to test reasoning across sources and longer time horizons. To address this gap, we present PROBE (Proactive Resolution of Bottlenecks). PROBE decomposes proactivity as a pipeline of three core capabilities: (1) searching for unspecified issues, (2) identifying specific bottlenecks, and (3) executing appropriate resolutions. We apply PROBE to evaluate leading LLMs and popular agentic frameworks, showing that even state-of-the-art models struggle to solve this benchmark. Computing our consistent measurements across frontier LLMs and agents, we find that the best end-to-end performance of 40% is achieved by both GPT-5 and Claude Opus-4.1. Additionally, we demonstrate the relative capabilities of each model and analyze mutual failure modes. Our results highlight the current limitations of autonomous action in agentic systems, and expose promising future research directions  [^1] Code and dataset available at https://github.com/fastino-ai/PROBE_benchmark.
@@ 1 Introduction
Figure 1: An end to end depiction of the PROBE task setup. The model (or agent) needs to use the world model to (i) search over a user datastore, (ii) identify the bottleneck and finally (iii) select the action to be executed. The model is evaluated across all tasks in this pipeline.
Prior studies on proactive agents have explored agent proactivity in interacting with physical environments (Zhang et al., 2023), asking follow-up questions (Zhang et al., 2024) and perceiving immediate needs from a personalized environment Lu et al. (2024a); Yang et al. (2025a). However, existing approaches compress evaluation into narrow, immediate temporal context, failing to capture insights that emerge only through longer-term analysis. For instance, proactive agents that look only at current context would not detect and take an appropriate action for a missed deadline from the past (as shown in figure 1). To this end, we operationalize proactivity as a three part construct. Given a set of priorities and a personalized user datastore, agents search across documents for user-relevant issues, identify the most pertinent ones (which we term bottlenecks), and resolve said issues by executing appropriate actions.
Constructing a real-world benchmark for proactivity is difficult since collecting long time horizon, multi-document user data raises privacy concerns and creates significant annotation overhead. Building on previous successes in the generation of synthetic datasets Nadas et al. (2025); Long et al. (2024); Butt et al. (2024), we construct a data generation agent to build our benchmark (we describe this in section 2). The resulting PROBE benchmark comprises of 1,000 diverse samples that challenge AI systems to proactively identify and resolve critical bottlenecks hidden within realistic workplace datastores (see figure1). Our evaluation reveals a striking capability gap: even state-of-the-art LLMs and specialized agentic frameworks achieve no more than 40% success on this end-to-end task, highlighting the substantial challenges that remain in developing truly proactive AI systems. From a model perspective, PROBE establishes a joint evaluation protocol for LLMs and agents, as shown in figure 1. In summary, our contribution in this paper is threefold:
  • We introduce PROBE - 1000 test samples that systematically evaluate proactive capabilities in AI systems through a unified framework, addressing a critical need for a realistic proactive benchmark.
@@ 2 Methodology
We develop a data generation pipeline to orchestrate our end-to-end workflow (described in figure 2). Starting from real user personas, we build comprehensive world models that capture each simulated user’s environment, goals, and constraints. These world models drive the creation of a datastore filled with personalized synthetic documents. While generating these documents, we ensure that several contain a prespecified hidden obstacle (bottleneck). For each bottleneck, our pipeline generates several candidate actions, only one of which resolves the issue. This setup requires agents to demonstrate genuine proactivity: they must discover the bottleneck through exploration and identify the right fix among several plausible options. The following sections detail our problem formulation and pipeline components.
@@ 2.1 Problem Definition
Consider a world-model $W$ of a user, let $D$ be a finite “universal” set of documents that constitutes the user’s datastore, which is a collection of all of user’s accessible documents, and let $b\in B$ denote a fixed bottleneck. We define a bottleneck as an issue that is critically important to the individual, actionable, and identifiable through a finite set of documents. For a given $b$, the rest of the documents that do not pertain to this bottleneck are considered distractors with respect to the bottleneck.
@@ 2.3 Data Generation Setup
Figure 2: End-to-end proactive benchmark generation pipeline. A synthetic world model ($W$) is constructed in reference to a LinkedIn user profile, followed by generating a bottleneck ($b$) to resolve. True Positives ($T\subseteq D$) and distractors (other documents unrelated to the bottleneck, $K=D\setminus T$) are then constructed to frame the bottleneck prediction task. Finally, a task to execute is selected with parameters ( $\mathcal{P}=\{P_{a}\}_{a\in A}$ for action $a\in A$ ) to resolve the bottleneck. For any bottleneck, only one of the candidate tasks is a plausible resolution.
We design a data generation pipeline that scales robustly in both context size and difficulty. Starting from real-world professional profiles, it constructs comprehensive synthetic world-models that mirror realistic workplace scenarios. Figure 2 illustrates the complete pipeline, which comprises four key components222all prompts for individual data generation modules are shown in appendix B:
@@
User Datastore: For each sample $S$, we construct the document set $D=T\cup K$. The True positives $T$ - documents where $f(d)=1$ - collectively provide sufficient evidence to identify bottleneck $b$. Distractors $K$ are documents where $f(d)=0$, introducing realistic noise with respect to the bottleneck. In our current datastore setup, all the generated documents are either emails, calendar events, or text documents, as exemplified in Figures  1 and  2.
To mirror real-world complexity, we employ two key design principles: (i) Evidence distribution: We often distribute evidence for $b$ across multiple documents in $T$, requiring agents to synthesize information from $t$ different sources. (ii) Contextual noise: We generate $k$ distractor documents of comparable length and professional relevance, ensuring bottleneck identification requires careful analysis rather than superficial pattern matching.
@@ 2.4 PROBE - benchmark for proactivity evaluation
We use the setup outlined above to generate the final dataset using GPT-4.1 as our primary model for generation. We generate a total of 1000 datapoints generated from 235 unique personas (described in 2.3). The full dataset stats are shown in Table 1333token counts measured using tiktoken https://platform.openai.com/tokenizer.
@@
To ensure the quality of the dataset, we adapt a modified approach from Jiang et al. (2025) by performing multiple rounds of filtering via an adversarial agent (GPT-5) on a small sample set (of 5 datapoints). After each round of data generation, we asked the adversarial agent to identify and exploit patterns in the data, preferring artifact-based solutions over independent reasoning. The final set of datapoints were generated only once no sample from our pipeline was solvable with these artifacts alone.
@@
To establish human performance benchmarks and validate task difficulty, we conducted a 4-hour annotation study with three annotators, all holding at least a master’s degree. Each annotator received identical instructions to those given to LLM systems and completed as many samples as possible within the specified time limit. Additionally, the annotators were instructed to judge if samples were realistic and feasible. To assess this, we asked the following questions: (i) Were the all documents you read realistic documents that you may see in a real workplace? and (ii) Were the actions you read realistic actions that you could see being used to resolve bottlenecks in a real workplace setting?
Annotator Search Bottleneck Ident Task Selection Entries Samples Realistic
ID F1 % -tification Score (%) Score (%) Annotated hour Artifacts
1 26.79 0.00 0.00 13 3.25
2 45.24 0.00 14.90 7 1.75
3 20.37 0.00 8.33 6 1.50
Avg 30.28 0.00 5.93 8.67 2.17
Table 2: Annotation numbers showing that humans could not accomplish our task successfully. Annotators on average retrieved 30%, never identified the bottleneck, and selected the correct task at just above random chance, labeling about 2 samples per hour.
Across 12 total annotator-hours, only 26 samples were successfully completed at an average throughput of 2.17 samples per hour per annotator. The annotation experiment showed the substantial cognitive load required for bottleneck identification across multiple documents, and the time-intensive nature of synthesizing evidence and selecting appropriate actions [^4] Annotators found no sample to be unrealistic, and gave feedback that the task was very challenging. For the annotators who answered “yes" to the two questions on judging artifacts, we record a green check mark in the “Realistic Artifacts" column of our results table. The annotation results for all three annotators shown in table 2. All human annotations were evaluated using the same metrics setup as described in section 3.
@@ 3.1 Metrics
Search: This evaluation measures how well an agent retrieves the relevant documents that are required for identifying the bottleneck precisely. We measure the agent’s retrieved document $\widehat{T}$ against the gold bottlenecks $T$ using standard precision, recall and F1 metrics.
@@
We provide the prompts for both the llm-as-a-judge prompts in appendix B.6.
@@ 3.2 LLM-as-a-judge
To validate our use of LLM-as-a-judge, we conducted a measurement study with 50 GPT-4.1 prediction-output pairs. Two human annotators independently evaluated bottleneck identification and parameter judgments for each sample. We achieved 84% inter-annotator agreement and 80% human-LLM agreement across 100 total annotations, supporting our decision to use LLM-based scoring.
@@ 3.3 Baselines
Models: We evaluate our benchmark against several frontier closed source models including OpenAI GPT-5(OpenAI, 2025), GPT-5-mini(OpenAI, 2025c), GPT-4.1(OpenAI, 2025a), GPT-4.1-mini(OpenAI, 2025b), Claude 4.1 Opus(Anthropic, 2025b), Claude 4 Sonnet(Anthropic, 2025a) and the best-performing open-source models including Kimi-K2(Team et al., 2025) and DeepSeek- R1(DeepSeek-AI et al., 2025). Among other open-source models, we test OpenAI GPT OSS(OpenAI, 2025) at both 120B and 20B scales [^5] We could not test Gemini-2.5 due to rate-limiting issues.
Agentic Frameworks: We evaluate three leading agentic frameworks: ReACT (Yao et al., 2023), Reflexion-Agent (Shinn et al., 2023), and ReWOO (Xu et al., 2024). Since these frameworks target different problem domains, we adapted each for bottleneck resolution. Our modifications include minor prompt changes, structured outputs, and two retrieval tools: embedding based search and SQL queries. Each framework takes a distinct approach. ReACT cycles between reasoning and retrieval, progressively building context until it converges on an action. Reflexion learns from its failures: it runs multiple trials of document retrieval, analysis, and action selection, using LLM-based reflection to improve after each unsuccessful attempt. ReWOO, in contrast, precoordinates the process. After constructing a structured plan, it dispatches specialized workers for search and reasoning tasks, then synthesizes their findings to pinpoint bottlenecks and select interventions. We did not include any pre-existing proactive agent frameworks (Lu et al., 2024b; Yang et al., 2025b) as baselines, as current systems are designed for specialized domains (conversational agents, UI navigation, embodied robotics). Resultantly, the distinct input modalities and task objectives of these frameworks do not trivially transfer to our workflow. All agentic frameworks use GPT-5-mini as the underlying base model. We provide more details in appendix A.
@@
Table 3: Comparative results across several frontier closed and open source models. GPT-5 and Claude Opus-4.1 show the best performance compared to the rest. We also show evaluation across search, bottleneck identification and task execution to show the relative strengths of weaknesses across models. Note: We found GPT-5-mini performance to be close to GPT-5 performance across the board, hence its results are removed for brevity.
@@ 3.4 Model Comparisons
Retrieval remains challenging: Our analysis reveals a clear performance hierarchy among models. Frontier models (GPT and Claude series) significantly outperform others (Kimi K-2, DeepSeek-R1, GPT-OSS series), with all models showing higher precision than recall - for example, GPT-5 achieves 0.73 precision but only 0.59 recall. This pattern suggests models retrieve conservatively, struggling to some extent to retrieve all relevant pieces of information needed to identify the bottleneck. This is especially pronounced in smaller Language Models, who seem to massively under-retrieve relevant documents and resultantly struggle with any sort of task selection.
@@ 3.5 Agent Frameworks Comparison
Even when agents generated plausible reasoning steps, the limited evidence prevented accurate diagnosis and parameter-complete actions. These results partly reflect the mismatch between our experimental constraints and typical agentic frameworks, which usually leverage web search, APIs, and environment interaction. Our benchmark’s restriction to SQL and semantic search limits the “act" channel that many agent architectures rely on for iterative knowledge acquisition.
@@ 4 Error Analysis
Incorrect root cause66footnotemark: 6 64.6% 70.6% 76.8% 72.1% 84.8%
@@
Action Selection and Parameter Prediction for actions : Action selection and selection of parameters for the action remains independently challenging. GPT-5 and Claude Opus achieve better parameter coverage but higher error rates (45.6% and 45.4% incorrect), while GPT-4.1 and Claude Sonnet show the inverse pattern; more accurate specification (35.3% and 36.1% incorrect) but higher miss rates. Current models cannot simultaneously achieve higher coverage and precise parameter specification within the complex workplace scenarios generated seen in this benchmark.
5 Ablation Studies
5.1 Impact of Context Window Size
5.2 Data Generation Model Diversity
6 Related Work
Emerging proactive agents face the challenge of anticipating user goals from partial observations and executing actions without instruction. Recent work explores intent inference from behavioral patterns (Zhao et al., 2025), continuous insight generation from data streams (Yang et al., 2025b), and context-aware action generation (Shaikh et al., 2025). However, these systems lack systematic evaluation frameworks that assess end-to-end proactive capabilities.
Agent Benchmarking: Existing benchmarks predominantly evaluate reactive systems: SWE-bench (Jimenez et al., 2024) for software engineering; ToolBench (Qin et al., 2023) for API calling; and GAIA (Mialon et al., 2023) for multi-hop reasoning. Recent proactivity benchmarks like ProactiveVideoQA (Wang et al., 2025) and ProCIS (Samarinas & Zamani, 2024) begin addressing this gap but remain limited to conversational actions. None of the above-mentioned works decompose proactivity into constituent capabilities or test comprehensive task execution across extended contexts and time horizons, motivating the systematic approach found in PROBE .

### v3

@@ Abstract
LLM-based agents are increasingly moving towards proactivity: rather than awaiting instruction, they exercise agency to anticipate user needs and solve them autonomously. However, evaluating proactivity is challenging; current benchmarks are constrained to localized context, limiting their ability to test reasoning across sources and longer time horizons. To address this gap, we present PROBE (Proactive Resolution of Bottlenecks). PROBE decomposes proactivity as a pipeline of three core capabilities: (1) searching for unspecified issues, (2) identifying specific bottlenecks, and (3) executing appropriate resolutions. We apply PROBE to evaluate leading LLMs and popular agentic frameworks, showing that even state-of-the-art models struggle to solve this benchmark. Computing our consistent measurements across frontier LLMs and agents, we find that the best end-to-end performance of 40% is achieved by both GPT-5 and Claude Opus-4.1. Additionally, we demonstrate the relative capabilities of each model and analyze mutual failure modes. Our results highlight the current limitations of autonomous action in agentic systems, and expose promising future research directions.
@@ 1 Introduction
Figure 1: An end-to-end depiction of the PROBE task setup. The model (or agent) needs to use the world model to (i) search over a user datastore, (ii) identify the bottleneck and finally (iii) select the action to be executed. The model is evaluated across all tasks in this pipeline.
Prior studies on proactive agents have explored agent proactivity in interacting with physical environments (Zhang et al., 2023), asking follow-up questions (Zhang et al., 2024) and perceiving immediate needs from a personalized environment Lu2024ProactiveAS; Yang et al. (2025b). However, existing approaches compress evaluation into narrow, immediate temporal context, failing to capture insights that emerge only through longer-term analysis. For instance, proactive agents that look only at current context would not detect and take an appropriate action for a missed deadline from the past (as shown in Figure 1). To this end, we operationalize proactivity as a three-part construct. Given a set of priorities and a personalized user datastore, agents search across documents for user-relevant issues, identify the most pertinent ones (which we term bottlenecks), and resolve said issues by executing appropriate actions.
Constructing a real-world benchmark for proactivity is difficult since collecting long time horizon, multi-document user data raises privacy concerns and creates significant annotation overhead. Building on previous successes in the generation of synthetic datasets Nadas et al. (2025); Long et al. (2024); Butt et al. (2024), we construct a data generation agent to build our benchmark (we describe this in section 2). The resulting PROBE benchmark comprises 1,000 diverse samples that challenge AI systems to proactively identify and resolve critical bottlenecks hidden within realistic workplace datastores (see Figure1). Our evaluation reveals a striking capability gap: even state-of-the-art LLMs and specialized agentic frameworks achieve no more than 40% success on this end-to-end task, highlighting the substantial challenges that remain in developing truly proactive AI systems. From a model perspective, PROBE establishes a joint evaluation protocol for LLMs and agents, as shown in Figure 1. In summary, our contribution in this paper is threefold:
  • We introduce PROBE - 1000 test samples that systematically evaluate proactive capabilities in AI systems through a unified framework, addressing a critical need for a long-horizon proactivity benchmark.
@@ 2 Methodology
We develop a data generation pipeline to orchestrate our end-to-end workflow (described in Figure 2). Starting from real user personas, we build comprehensive world models that capture each simulated user’s environment, goals, and constraints. These world models drive the creation of a datastore filled with personalized synthetic documents. While generating these documents, we ensure that several contain a pre-specified hidden obstacle (bottleneck). For each bottleneck, our pipeline generates several candidate actions, only one of which resolves the issue. This setup requires agents to demonstrate genuine proactivity: they must discover the bottleneck through exploration and identify the right fix among several plausible options. The following sections detail our problem formulation and pipeline components.
@@ 2.1 Problem Definition
Consider a world-model $W$ of a user, let $D$ be a finite “universal” set of documents that constitutes the user’s datastore, which is a collection of all of the user’s accessible documents, and let $b\in B$ denote a fixed bottleneck. We define a bottleneck as an issue that is critically important to the individual, actionable, and identifiable through a finite set of documents. For a given $b$, the rest of the documents that do not pertain to this bottleneck are considered distractors with respect to the bottleneck.
@@ 2.3 Data Generation Setup
Figure 2: End-to-end benchmark generation pipeline. A synthetic world model ($W$) is constructed from a real LinkedIn profile, along with a bottleneck ($b$) to resolve. True Positives documents containing the bottleneck ($T\subseteq D$) and irrelevant Distractors $(K=D\setminus T$) are then generated: the agent will search through both to predict the bottleneck. Finally, an "available action" ( $\mathcal{P}=\{P_{a}\}_{a\in A}$) is selected to resolve the bottleneck.
We design a data generation pipeline that scales robustly in both context size and difficulty. Starting from real-world professional profiles, it constructs comprehensive synthetic world-models that mirror realistic workplace scenarios. Figure 2 illustrates the complete pipeline, which comprises four key components111all prompts for individual data generation modules are shown in Appendix D:
@@
User Datastore: For each sample $S$, we construct the document set $D=T\cup K$. The True positives $T$ - documents where $f(d)=1$ - collectively provide sufficient evidence to identify bottleneck $b$. Distractors $K$ are documents where $f(d)=0$, introducing realistic noise with respect to the bottleneck. In our current datastore setup, all the generated documents are typical emails, calendar events, or text documents, as exemplified in Figures  1 and  2. All generated documents are personalized, as generations are contextualized by the LinkedIn persona and world model.
To mirror real-world complexity, we employ two key design principles: (i) Evidence distribution: We often distribute evidence for $b$ across multiple documents in $T$, requiring agents to synthesize information from $t$ different sources. These documents are relevantly timestamped, meaning the task often requires both contextual and temporal reasoning. (ii) Contextual noise: We generate $k$ distractor documents of comparable length and professional relevance, ensuring bottleneck identification requires careful analysis rather than superficial pattern matching.
@@ 2.4 PROBE - benchmark for proactivity evaluation
We use the setup outlined above to generate the final dataset using GPT-4.1 as our primary model for generation. We generate a total of 1000 data points generated from 235 unique personas (described in 2.3).
@@
The full dataset stats are shown in Table 1222token counts measured using tiktoken https://platform.openai.com/tokenizer. We also provide a listing of bottleneck categories in Figure 3, as well as a complete world model-bottleneck combination in Appendix C.
To ensure the faithfulness of the dataset, we adapt a modified approach from Jiang et al. (2025) by performing multiple rounds of filtering via an adversarial agent (GPT-5) on a small sample set (of 5 data points). After each round of data generation, we asked the adversarial agent to identify and exploit patterns in the data, preferring artifact-based solutions over independent reasoning. The final set of data points was generated only once no sample from our pipeline was solvable with these artifacts alone.
Figure 3: Distribution of bottleneck categories in the PROBE benchmark. The 1,000 synthetic test instances span six distinct workplace scenarios, ranging from pending deliverables and reviews to training gaps and communication breakdowns. These categories reflect realistic organizational challenges that proactive agents must identify and resolve through multi-document reasoning.
@@
To establish human performance benchmarks and validate task difficulty, we conducted an annotation study with three annotators, all holding math or computer science degrees. Each annotator received identical instructions to those given to LLM systems and completed as many samples as possible within the specified time limit. Though the original context length of the benchmark (100k tokens) was far too long to allow for reasonable human performance or collecting a meaningful number of attempts, we were able to validate the identification and task execution components using a "Short-Context" set. We did this by by removing distractors and summarizing true positives, thus reducing the input to $\approx$ 1,000 tokens.
Annotator Identification Task Execution Samples
ID Score Score Annotated
1 0.86 0.88 25
2 0.46 0.56 25
3 0.82 0.56 25
Average 0.71 0.67 Total: 75
Table 2: Human evaluation results after removing distractors and summarizing true positives. The substantial improvement in annotator performance indicates that the main challenge of the task stems from context length.
As shown in Table 2, humans perform reasonably in these conditions. We also record substantial inter-annotator agreement (Fleiss’ $\kappa=0.714$), providing evidence that the task is well defined: the difficulty stems significantly from the challenge of long-context reasoning.
@@ 3.1 Metrics
Search: This evaluation measures how well an agent retrieves the relevant documents that are required for identifying the bottleneck precisely. We measure the agent’s retrieved documents $\widehat{T}$ against the gold bottlenecks $T$ using standard precision, recall and F1 metrics.
@@
We provide both LLM-as-a-judge prompts in Appendix  D.6.
@@ 3.2 LLM-as-a-judge
To validate our use of LLM-as-a-judge, we conducted a measurement study with 100 uniformly sampled GPT-4.1 prediction-output pairs. Two human annotators independently evaluated bottleneck identification and parameter judgments for each sample. We achieved 85% inter-annotator agreement (Fleiss’ $\kappa=0.7$: substantial) and 81% (Fleiss’ $\kappa=0.62$: substantial) human-LLM agreement across 100 total annotations, supporting our decision to use LLM-based scoring.
@@ 3.3 Baselines
Models: We evaluate our benchmark against several frontier closed source models including OpenAI GPT-5 (OpenAI, 2025b), GPT-5-mini (OpenAI, 2025e), GPT-4.1 (OpenAI, 2025d), GPT-4.1-mini (OpenAI, 2025a), Claude 4.1 Opus (Anthropic, 2025a), Claude 4 Sonnet (Anthropic, 2025b) and the best-performing open-source models including Kimi-K2 (Team et al., 2025) and DeepSeek- R1 (DeepSeek-AI et al., 2025). Among other open-source models, we test OpenAI GPT OSS (OpenAI, 2025c) at both 120B and 20B scales [^3] We could not test Gemini-2.5 due to rate-limiting issues.
Agentic Frameworks: We evaluate three leading agentic frameworks: ReACT (Yao et al., 2023), Reflexion-Agent (Shinn et al., 2023), and ReWOO (Xu et al., 2024). Since these frameworks target different problem domains, we adapted each for bottleneck resolution. Our modifications include minor prompt changes, structured outputs, and two retrieval tools: embedding-based search and SQL queries. Each framework takes a distinct approach. ReACT cycles between reasoning and retrieval, progressively building context until it converges on an action. Reflexion learns from its failures: it runs multiple trials of document retrieval, analysis, and action selection, using LLM-based reflection to improve after each unsuccessful attempt. ReWOO, in contrast, pre-coordinates the process. After constructing a structured plan, it dispatches specialized workers for search and reasoning tasks, then synthesizes their findings to pinpoint bottlenecks and select interventions. We did not include any pre-existing proactive agent frameworks (Lu et al., 2024; Yang et al., 2025a) as baselines, as current systems are designed for specialized domains (conversational agents, UI navigation, embodied robotics). Resultantly, the distinct input modalities and task objectives of these frameworks do not trivially transfer to our workflow. All agentic frameworks use GPT-5-mini as the underlying base model. We provide more details in Appendix B.
@@
Table 3: Comparative results across several frontier closed and open source models. GPT-5 and Claude Opus-4.1 show the best performance compared to the rest. We also show evaluation across search, bottleneck identification and task execution to show the relative strengths and weaknesses across models. Note: We found GPT-5-mini performance to be close to GPT-5 performance across the board, hence its results are removed for brevity.
@@ 3.5 Agent Frameworks Comparison
These results should be interpreted in context of the task structure. Standard agentic frameworks typically perform iterative, open-ended search over large external tools (e.g., web search, APIs). Our proactive task, however, requires agents to explore an unfamiliar datastore without a well-defined search target at the outset. This negates the benefit these frameworks attain from external tool use. Notably, all agentic frameworks are equitably evaluated: all access the data through SQL and Semantic Search.
@@ 4 Error Analysis
Incorrect root cause44footnotemark: 4 64.6% 70.6% 76.8% 72.1% 84.8%
@@
Action Selection and Parameter Prediction for actions : Action selection and selection of parameters for the action remains independently challenging. GPT-5 and Claude Opus achieve better parameter coverage but higher error rates (45.6% and 45.4% incorrect), while GPT-4.1 and Claude Sonnet show the inverse pattern; more accurate specification (35.3% and 36.1% incorrect) but higher miss rates. Current models cannot simultaneously achieve higher coverage and precise parameter specification within the complex workplace scenarios seen in this benchmark.